OpenStudy (anonymous):

What is the denominator of the simplified form of (1/2) - (3 / 2x + 1) 2 2x + 1 2(2x + 1) 4

OpenStudy (anonymous):

2x+1 is the answer

OpenStudy (anonymous):

2(2x+1)

OpenStudy (anonymous):

I thought it was 2(2x + 1) too

OpenStudy (anonymous):

That is the wrong answer

OpenStudy (anonymous):

Since you can factor the numerator and then divide the 2

OpenStudy (anonymous):

it comes to 2x - 5 ----- 2(2x+1)

OpenStudy (anonymous):

The numerator is 2x-2

OpenStudy (anonymous):

since 2x+1-3=2x-2

OpenStudy (anonymous):

And then factor the numerator as 2(x-1)

OpenStudy (anonymous):

and then divid the 2s which results in (x-1)/(2x+1)

OpenStudy (anonymous):

Do u get it?

OpenStudy (anonymous):

My ending result was (2x - 5) / 2 (2x + 1)

OpenStudy (anonymous):

Why did you get 2x-5?

OpenStudy (anonymous):

2x+1 -3=2x-2

OpenStudy (anonymous):

Wait How did you get 2x - 2

OpenStudy (anonymous):

1(2x+1)=2x+1 and -3(1)=-3 2x+1-3=2x-2

OpenStudy (anonymous):

mitosuki - is the denominator of the last term 2x+1 or is it (3/2x ) + 1 ?

OpenStudy (anonymous):

it is 2x + 1

OpenStudy (anonymous):

3 _____ 2x + 1 It looks like that

OpenStudy (anonymous):

then you are correct

OpenStudy (anonymous):

When you are trying to get the same denomanators u are most probably doing something wrong

OpenStudy (anonymous):

lol - im going crazy here!! the lcm of the two fractions is 2(2x+1) - agreed?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Yes, that's what I got

OpenStudy (anonymous):

To the first fraction you multiply the denomanator and numerator by (2x+1) and the second you multiply the numerator and denomanator by 2

OpenStudy (anonymous):

I got something wrong!!!!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Jimmy you were correct and mitosuki

OpenStudy (anonymous):

I am just giving u a run for your money!!!!

OpenStudy (anonymous):

lol - ok i believe you!!

OpenStudy (anonymous):

I forgot to multiply the numerator of the second fraction by 2

OpenStudy (anonymous):

And guess what I taught simular material in class today!!!

OpenStudy (anonymous):

LOL Thank you guys :D